OpenStudy (anonymous):

Harold is graphing the solution to the equation x + 4y = 2. He begins by making the table below. What is the missing number?

OpenStudy (anonymous):



OpenStudy (anonymous):

I believe you would set this up in the slope intercept form and then jsut plug in the value of x. Right?

OpenStudy (anonymous):

Just plug in the x's and see what the outcome it?

OpenStudy (anonymous):

* is

OpenStudy (anonymous):

Well if you plug in the value of x the the equation should equal the value of y. So here is an example, using the numbers in the first column. x+4y=2 4y=-x+2 y=-1/4+.5 Then you plug in the values y=-1/4(-4)+.5 Do you understand what I did to get that answer?

OpenStudy (anonymous):

oops sorry 3/2=-1/4(-4)+.5

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok then you know how to do it :) Hopefully I explained it well enough.

OpenStudy (anonymous):

When i solved it I got 3/2 = 3/2 ?

OpenStudy (anonymous):

oh, then do I plug in 0 next?

OpenStudy (anonymous):

Are oyu talking about the example I just did from the given numbers?

OpenStudy (anonymous):

yes. but to get the missing number do I plug in 0 into the equation?

OpenStudy (anonymous):

Yeah.

OpenStudy (anonymous):

i got 1/2

OpenStudy (anonymous):

So 1/2 is the missing number. right?

OpenStudy (anonymous):

Yup. Thats what I got too. :)